ganeshie8 (ganeshie8):

Find the number of values of \(n\) between \(1\) and \(100\) such that \(x^2+x-n=0\) has integer roots.

OpenStudy (fibonaccichick666):

there's only 2?

ganeshie8 (ganeshie8):

not sure im still working on this haha! but my first guess would be more than 2

OpenStudy (fibonaccichick666):

I mean only two for each n with consecutive factor pairs

OpenStudy (fibonaccichick666):

so I would expect there to be 9 total

OpenStudy (fibonaccichick666):

pairs

OpenStudy (fibonaccichick666):

with non distinct values

ganeshie8 (ganeshie8):

x^2+x-2 = (x-1)(x+2) how did u get 9 pairs so fast ! xD

OpenStudy (agent0smith):

n and n+1, where 1<n<9?

OpenStudy (fibonaccichick666):

well, they have to be consecutive to give us a positive 1 in the middle for the coefficient right? they cannot multiply to more than 100, so since 10*11 >100, we only need pairs less than 10

ganeshie8 (ganeshie8):

Brilliant !!!

OpenStudy (fibonaccichick666):

(-1,2)=-2 (-2,3)=-6 (-3,4)=-12 yada yda till we hit, (-9,10)=-90

OpenStudy (agent0smith):

I should've used a diff variable, not n. But same thing.

OpenStudy (fibonaccichick666):

so the limit on n makes our life easier

ganeshie8 (ganeshie8):

I was trying quadratic formula and making it hard

OpenStudy (fibonaccichick666):

oh, lol. I was about to say, this is an easy one for your genius mind

ganeshie8 (ganeshie8):

\[x = \dfrac{-1\pm \sqrt{1+4n}}{2}\] \(1+4n = m^2\) etc..

OpenStudy (agent0smith):

Well they have to be integers, so that makes it easy

ganeshie8 (ganeshie8):

my start was wrong.. I see it is easy with pairs idea :)

OpenStudy (fibonaccichick666):

yea, no forest for trees :)

OpenStudy (fibonaccichick666):

I'm just excited because I actually answered a question you asked!

ganeshie8 (ganeshie8):

this is the fastest somebody ever helped me haha xD

OpenStudy (fibonaccichick666):

lol

OpenStudy (agent0smith):

The question looks a lot more complex when Openstudy is being weird...

ganeshie8 (ganeshie8):

omg! that happened to me yesterday too... i think it is some latex related bug

OpenStudy (fibonaccichick666):

... that problem would be insane,

OpenStudy (fibonaccichick666):

just try it with 111 and 100,100,100 for your n.... we would have 12-> like somewhere between10,000 and 11,000 for our restraints... That's alot of pairs

OpenStudy (agent0smith):

I'm pretty sure @ganeshie8's sole reason for asking this question, was so that he would have equal numbers of fan testimonials as questions asked.

OpenStudy (fibonaccichick666):

lol

ganeshie8 (ganeshie8):

lol that would be just a coincidence XD Fib the answer would be number of perfect squares between 111 and 100,100,100 right ?

ganeshie8 (ganeshie8):

(from quadratic formula)

OpenStudy (fibonaccichick666):

no,

OpenStudy (fibonaccichick666):

it would be the number of consecutive pairs within \(12\le x,y\le 10005\)

ganeshie8 (ganeshie8):

Ahh okay, i think it should be number of ODD perfect squares between `111` and `1+4*100,100,100`

OpenStudy (fibonaccichick666):

that should be something like 10005-12=9993

OpenStudy (fibonaccichick666):

by my method

ganeshie8 (ganeshie8):

YES! that kinda checks out with my method too number of perfect squares less than `1+4*100,100,100` would be `floor(sqrt(1+4*100,100,100))` = 20010 subtract perfect squares less than 111 and divide by 2

OpenStudy (fibonaccichick666):

I don't know how to find all of those perfect squares though

ganeshie8 (ganeshie8):

ikr :) question is about the count, not the values of n :P

OpenStudy (fibonaccichick666):

true haha, but we did if your method works, just find a way to determine how many perfect squares are between two numbers

ganeshie8 (ganeshie8):

number of perfect squares between \(a\) and \(b\) would be somewhere around \(\sqrt{b} - \sqrt{a}\) right ?

ganeshie8 (ganeshie8):

assuming \(a < b\)

OpenStudy (fibonaccichick666):

uhm, that I do not know, shall we attempt that?

OpenStudy (fibonaccichick666):

see if we get close to my and your answer

ganeshie8 (ganeshie8):

idk thats just a ballpark number....

OpenStudy (fibonaccichick666):

mine should be exact

ganeshie8 (ganeshie8):

do we agree number of perfect squares not greater than \(n\) is \(\lfloor \sqrt{n}\rfloor \)

ganeshie8 (ganeshie8):

exampel : number of perfect squares not greater than \(256\) is \(\lfloor \sqrt{256}\rfloor = 16 \) : \[\{1^2,2^2,3^2,\ldots, 16^2\}\]

OpenStudy (fibonaccichick666):

9994.46809437 Pretty dang close

OpenStudy (fibonaccichick666):

I don't know the floor function

ganeshie8 (ganeshie8):

floor(2.9) = 2 floor(2) = 2 floor(1.9) = 1 floor(n) is the greatest integer less than or eqal to n

OpenStudy (fibonaccichick666):

ohk, so where did you get that theorem about perfect squares from?

ganeshie8 (ganeshie8):

observation

OpenStudy (fibonaccichick666):

hmmm, I think we may have just derived the phi function or something like that from number theory, but I'm tired, and thus, cannot be bothered to check if that is the case

ganeshie8 (ganeshie8):

thats okay but it has nothing to do with phi function

OpenStudy (fibonaccichick666):

it's one of those number theory thingies i think